Geologists use the decay of potassium-40 in volcanic rocks to determine their age. Potassium-40 has a half-life of 1.26 x 10⁹ years, so it can be used to date very old rocks. If a sample of rock 3.78 x 10⁹ years old contains 2.73 x 10⁻⁷ g of potassium-40 today, how much potassium-40 was originally present in the rock? *
1
Expert's answer
2019-10-17T07:31:37-0400

The exponential decay of potassium-40 can be described by the following equation:

Nt = N0 × e-λt,

where Nt - the remaining quantity of atoms, N0 - the initial quantity, λ - the decay constant (λ = 0.693 / t1/2), t - the mean lifetime, t1/2 - the half-life.

From here:

N0 = Nt / e-λt.

Nt = NA × (mK / MrK),

where NA - Avogadro's number, mK - potassium-40 mass, MrK - atomic weight.

As a result:

Nt = 6.022 × 1023 mol-1 × (2.73 × 10⁻⁷ g / 40 g/mol) = 0.41 × 1016

λ = 0.693 / 1.26 × 109 = 0.55 × 10-9.

Finally, the initial number of potassium-40 atoms:

N0 = 0.41 × 1016 / e-(0.55 × 10-9 × 3.78 x 10⁹) = 3.28 × 1016

From here:

m = Mr × N0/NA = 40 × (3.28 × 1016 / 6.022 × 1023) = 21.79 × 10-7 g.


Answer: 21.79 × 10-7 g of potassium-40 was originaly present in the rock.